What is the total impedance at 20kHz of a series circuit consisting of a 1.5mH inductance, a 100Ω resistance, and a 0.08μF capacitance?


1
Expert's answer
2022-01-31T10:13:56-0500

The total impedance of the circuit is


"Z=\sqrt{R^2+(2\pi f L-1/(2\pi f C))^2}=133.9\Omega."